# Efficient $\varepsilon$-approximate minimum-entropy couplings

### v1

@@ I-B Prior algorithmic guarantees
The first works on approximation guarantees studied a collection of algorithms with respect to a majorization lower bound. Cicalese, Gargano, and Vaccaro [22] introduced an algorithm that is 11-additive when $m=2$, and $\lceil\log(m)\rceil$-additive for general $m$: yielding the first constant-additive guarantee for constant $m$. The work of Kocaoglu, Dimakis, Vishwanath, and Hassibi [8] introduced the greedy coupling algorithm, that was later shown to be a local optimum by the same authors [10], and a 11-additive algorithm when $m=2$ by Rossi [24]. Later, Li [23] introduced a new algorithm that is $(2-2^{2-m})$-additive: yielding the first additive constant for general $m$. Compton [25] improved the greedy coupling guarantee to $\log(e)\approx 1.44$ for general $m$, while also showing a barrier in how the approximation analysis was tight with respect to the majorization lower bound.
@@ I-C Our result
For any $0<\varepsilon<\nicefrac{{1}}{{2}}$, there exists an algorithm with running time $(n/\varepsilon)^{O(m^{3}\cdot 2^{6m}\cdot\log^{3}(1/\varepsilon)/\varepsilon^{2})}$ that outputs a coupling $\operatorname{ALG}$, where $H(\operatorname{ALG})\leq H(\operatorname{OPT})+\varepsilon$.
@@ IV-A Preprocessing
Consider splitting a probability mass $x$ into two states cxcx and (1c)x(1-c)x for $c\leq\nicefrac{{1}}{{2}}$. The entropy is bounded by
@@
Consider splitting a probability mass $x$ into two states, where the smaller state is at most cxcx for $c\leq 1/e$. Then, the entropy is bounded by
@@
In this lemma, we focus on some input distribution $p_{i}$. For some distribution state with initial value $x\triangleq p_{i}(j)$, let us track the recursive rounding process with some vector, where $x_{0}=x$, then $x_{1}$ is the value of xyx-y in the first round, and so on. By Eq. 1, we know $x_{i+1}\leq\varepsilon^{2}x_{i}$. Let $\Delta_{i}$ denote the entropy increase to $\operatorname{SOL}^{\prime}$ from the split at phase ii. Suppose phase $i$ caused some coupling state of $\operatorname{SOL}^{\prime}$ with mass $c_{i}$ to be split into $d_{i}$ and $c_{i}-d_{i}$ (where we choose $d_{i}<c_{i}-d_{i}$). Then using Corollary IV.3:
@@
Note how the sum of the values of $x_{0}$ for each state of $p_{i}$ is 11. This implies that after all states are rounded for some distribution, the solution entropy will increase by at most $2\varepsilon^{2}\cdot(e+\log(1/\varepsilon^{2}))$. For $\varepsilon<\nicefrac{{1}}{{10}}$, this is at most $2\varepsilon$. ∎
@@ IV-C Matching
This gives us a clear modification plan. We will split the coupling state of size $q$ at the points $z$ and $z+a_{i}$ for all ii. By the previous statement, all $z+a_{i}\in[z,q)$. Observe how this modified $\operatorname{SOL}^{\prime}$ is now a valid coupling after each distribution state was split into $z,a_{i},b_{i}$: the coupling state of size $z$ will map to the new distribution state of size $z$, all coupling states ending by $z+a_{i}$ will map to the new distribution state of size $a_{i}$, and the remaining coupling states will be left with the distribution state of size $b_{i}$. In total, the qzq-z remaining coupling mass was split into at most $m+1$ new coupling states. In terms of bookkeeping, we will now ignore the new coupling state of size $z$ since it is fully coupled, all coupling states in the new distribution state of size $a_{i}$ will be considered regular, in the distribution state $b_{i}$ we will consider its new coupling states as regular (these are the coupling states that came from the range $z+a_{i},q]$), but the existing coupling states in $b_{i}$ will retain their previous type. This is depicted in Fig. 3. Informally, it may be helpful to think of this process as modifying a distribution state $\operatorname{Ind}_{i}$ by “deleting” $z$ mass, forking off $a_{i}$ mass into a new distribution state, and $b_{i}$ mass is what remains of the initial state $\operatorname{Ind}_{i}$.
@@
Since the sum of $z$ matched throughout the entire process is at most 11, these matching operations increase entropy by at most $\varepsilon\cdot(4\log(1/\varepsilon)+2\log(m+1))\leq 6\varepsilon\log(1/\varepsilon)$, since $1/\varepsilon\geq 2^{3m}\cdot 4m>m+1$.
@@ IV-D Splitting
Figure 5: Depiction of how a coupling state is modified over time. Nodes in this tree represent a coupling state, and outgoing edges represent when a coupling state is modified, resulting in its children. The root is a coupling state in $\operatorname{SOL}^{\prime}$ at the start of the dynamic program. Solid outgoing arrows represent a modification caused by a splitting operation (the sum of the two child sizes will be the same as the parent size), while dashed arrows represent a modification caused by a matching operation (the at most $m+1$ children will each have size at most $2\varepsilon$ fraction of its parent via $q-z\leq 2\varepsilon z\leq 2\varepsilon q$). The near-descendants of a node is the set of nodes that can be reached by a downward path of exclusively solid edges; e.g. the near-descendants of $a$ are $\{b,c,g,h\}$. The near-height of a node is the number of edges in the longest downward path from the node, consisting of only solid edges; the near-height of $a$ is 22. The level of a node is the number of dashed edges in the path from the root to this node; the level of $a$ is 0, the level of $k$ is 2.
@@
The total sum of $z$ that are matched is at most 11, and each $z$ increases the sum of $\operatorname{Matched}$ quantities by mzmz. ∎
@@
For each distribution, the total mass remaining is at most 11, implying this claim. ∎
@@ IV-E Leftover
Since both series have consecutive ratios of at least 22, using $4m\varepsilon\leq 1/4$, it holds:
@@ IV-F Concluding the analysis
for $C=3726m\cdot 2^{3m}$. Hence, if $\varepsilon^{\prime}\leq\frac{\varepsilon}{3726m\cdot 2^{3m}\cdot\log(1/\varepsilon)}$, then we achieve our desired approximation error and the value of $\varepsilon^{\prime}$ satisfies the required condition that $\varepsilon^{\prime}\leq\frac{1}{2^{3m}\cdot 4m}$. Since we additionally require that $\varepsilon^{\prime}$ is a power of 22, we simply round down to the nearest power of 22, and conclude this is a valid $\varepsilon^{\prime}$ where $\varepsilon^{\prime}\geq\frac{\varepsilon}{7452m\cdot 2^{3m}\cdot\log(1/\varepsilon)}$. Moreover, since $\alpha=\varepsilon^{3}/2$ and $\tau=\varepsilon^{9}/n^{6}$, we conclude the running time bound of
$(1/\tau)^{O(m\log(\nicefrac{{1}}{{(\varepsilon^{\prime}\alpha)}})/(\varepsilon^{\prime})^{2})}\leq(n/\varepsilon)^{O(m\cdot(\log(1/\varepsilon)+m)\cdot(\log^{2}(1/\varepsilon)\cdot m^{2}\cdot 2^{6m}/\varepsilon^{2}))}\leq(n/\varepsilon)^{O(m^{3}\cdot 2^{6m}\cdot\log^{3}(1/\varepsilon)/\varepsilon^{2})}.\qed$
@@ V Discussion
In Theorem I.1, we showed how there exists an $\varepsilon$-approximate coupling algorithm with running time $(n/\varepsilon)^{O(m^{3}\cdot 2^{6m}\cdot\log^{3}(1/\varepsilon)/\varepsilon^{2})}$: proving there exists a PTAS for constant $m=O(1)$. We suspect that techniques similar to those in this work could likely yield an analogous result for a similar problem studied in [37], although we do not execute this plan ourselves.

### v2

@@ I-B Prior algorithmic guarantees
The first works on approximation guarantees studied a collection of algorithms with respect to a majorization lower bound. Cicalese, Gargano, and Vaccaro [22] introduced an algorithm that is $1$-additive when $m=2$, and $\lceil\log(m)\rceil$-additive for general $m$: yielding the first constant-additive guarantee for constant $m$. The work of Kocaoglu, Dimakis, Vishwanath, and Hassibi [8] introduced the greedy coupling algorithm, that was later shown to be a local optimum by the same authors [10], and a $1$-additive algorithm when $m=2$ by Rossi [24]. Later, Li [23] introduced a new algorithm that is $(2-2^{2-m})$-additive: yielding the first additive constant for general $m$. Compton [25] improved the greedy coupling guarantee to $\log(e)\approx 1.44$ for general $m$, while also showing a barrier in how the approximation analysis was tight with respect to the majorization lower bound.
@@ I-C Our result
For any $0<\varepsilon<\nicefrac{{1}}{{2}}$, there exists an algorithm with running time $(n/\varepsilon)^{O(m^{4}\cdot 2^{6m}\cdot\log^{3}(1/\varepsilon)/\varepsilon^{2})}$ that outputs a coupling $\operatorname{ALG}$, where $H(\operatorname{ALG})\leq H(\operatorname{OPT})+\varepsilon$.
@@ IV-A Preprocessing
Consider splitting a probability mass $x$ into two states $cx$ and (1c)x(1-c)x for $c\leq\nicefrac{{1}}{{2}}$. The entropy is bounded by
@@
Consider splitting a probability mass $x$ into two states, where the smaller state is at most $cx$ for $c\leq 1/e$. Then, the entropy is bounded by
@@
In this lemma, we focus on some input distribution $p_{i}$. For some distribution state with initial value $x\triangleq p_{i}(j)$, let us track the recursive rounding process with some vector, where $x_{0}=x$, then $x_{1}$ is the value of xyx-y in the first round, and so on. By Eq. 1, we know $x_{i+1}\leq\varepsilon^{2}x_{i}$. Let $\Delta_{i}$ denote the entropy increase to $\operatorname{SOL}^{\prime}$ from the split at phase $i$. Suppose phase $i$ caused some coupling state of $\operatorname{SOL}^{\prime}$ with mass $c_{i}$ to be split into $d_{i}$ and $c_{i}-d_{i}$ (where we choose $d_{i}<c_{i}-d_{i}$). Then using Corollary IV.3:
@@
Note how the sum of the values of $x_{0}$ for each state of $p_{i}$ is $1$. This implies that after all states are rounded for some distribution, the solution entropy will increase by at most $2\varepsilon^{2}\cdot(e+\log(1/\varepsilon^{2}))$. For $\varepsilon<\nicefrac{{1}}{{10}}$, this is at most $2\varepsilon$. ∎
@@ IV-C Matching
This gives us a clear modification plan. We will split the coupling state of size $q$ at the points $z$ and $z+a_{i}$ for all $i$. By the previous statement, all $z+a_{i}\in[z,q)$. Observe how this modified $\operatorname{SOL}^{\prime}$ is now a valid coupling after each distribution state was split into $z,a_{i},b_{i}$: the coupling state of size $z$ will map to the new distribution state of size $z$, all coupling states ending by $z+a_{i}$ will map to the new distribution state of size $a_{i}$, and the remaining coupling states will be left with the distribution state of size $b_{i}$. In total, the qzq-z remaining coupling mass was split into at most $m+1$ new coupling states. In terms of bookkeeping, we will now ignore the new coupling state of size $z$ since it is fully coupled, all coupling states in the new distribution state of size $a_{i}$ will be considered regular, in the distribution state $b_{i}$ we will consider its new coupling states as regular (these are the coupling states that came from the range $z+a_{i},q]$), but the existing coupling states in $b_{i}$ will retain their previous type. This is depicted in Fig. 3. Informally, it may be helpful to think of this process as modifying a distribution state $\operatorname{Ind}_{i}$ by “deleting” $z$ mass, forking off $a_{i}$ mass into a new distribution state, and $b_{i}$ mass is what remains of the initial state $\operatorname{Ind}_{i}$.
@@
Since the sum of $z$ matched throughout the entire process is at most $1$, these matching operations increase entropy by at most $\varepsilon\cdot(4\log(1/\varepsilon)+2\log(m+1))\leq 6\varepsilon\log(1/\varepsilon)$, since $1/\varepsilon\geq 2^{3m}\cdot 4m>m+1$.
@@ IV-D Splitting
Figure 5: Depiction of how a coupling state is modified over time. Nodes in this tree represent a coupling state, and outgoing edges represent when a coupling state is modified, resulting in its children. The root is a coupling state in $\operatorname{SOL}^{\prime}$ at the start of the dynamic program. Solid outgoing arrows represent a modification caused by a splitting operation (the sum of the two child sizes will be the same as the parent size), while dashed arrows represent a modification caused by a matching operation (the at most $m+1$ children will each have size at most $2\varepsilon$ fraction of its parent via $q-z\leq 2\varepsilon z\leq 2\varepsilon q$). The near-descendants of a node is the set of nodes that can be reached by a downward path of exclusively solid edges; e.g. the near-descendants of $a$ are $\{b,c,g,h\}$. The near-height of a node is the number of edges in the longest downward path from the node, consisting of only solid edges; the near-height of $a$ is $2$. The level of a node is the number of dashed edges in the path from the root to this node; the level of $a$ is 0, the level of $k$ is 2.
@@
The total sum of $z$ that are matched is at most $1$, and each $z$ increases the sum of $\operatorname{Matched}$ quantities by $mz$. ∎
@@
For each distribution, the total mass remaining is at most $1$, implying this claim. ∎
@@ IV-E Leftover
Since both series have consecutive ratios of at least $2$, using $4m\varepsilon\leq 1/4$, it holds:
@@ IV-F Concluding the analysis
for $C=3726m\cdot 2^{3m}$. Hence, if $\varepsilon^{\prime}\leq\frac{\varepsilon}{3726m\cdot 2^{3m}\cdot\log(1/\varepsilon)}$, then we achieve our desired approximation error and the value of $\varepsilon^{\prime}$ satisfies the required condition that $\varepsilon^{\prime}\leq\frac{1}{2^{3m}\cdot 4m}$. Since we additionally require that $\varepsilon^{\prime}$ is a power of $2$, we simply round down to the nearest power of $2$, and conclude this is a valid $\varepsilon^{\prime}$ where $\varepsilon^{\prime}\geq\frac{\varepsilon}{7452m\cdot 2^{3m}\cdot\log(1/\varepsilon)}$. Moreover, since $\alpha=\varepsilon^{3}/2$ and $\tau=\varepsilon^{9}/n^{6}$, we conclude the running time bound of
$(1/\tau)^{O(m\log(\nicefrac{{1}}{{(\varepsilon^{\prime}\alpha)}})/(\varepsilon^{\prime})^{2})}\leq(n/\varepsilon)^{O(m\cdot(\log(1/\varepsilon)+m)\cdot(\log^{2}(1/\varepsilon)\cdot m^{2}\cdot 2^{6m}/\varepsilon^{2}))}\leq(n/\varepsilon)^{O(m^{4}\cdot 2^{6m}\cdot\log^{3}(1/\varepsilon)/\varepsilon^{2})}.\qed$
@@ V Discussion
In Theorem I.1, we showed how there exists an $\varepsilon$-approximate coupling algorithm with running time $(n/\varepsilon)^{O(m^{4}\cdot 2^{6m}\cdot\log^{3}(1/\varepsilon)/\varepsilon^{2})}$: proving there exists a PTAS for constant $m=O(1)$. We suspect that techniques similar to those in this work could likely yield an analogous result for a similar problem studied in [37], although we do not execute this plan ourselves.